In a National Achievement Test, suppose three test booklets are tested at random. Let D represent the defective test booklet and let N represent the non-defective test booklet. Let X be the random variable representing the number of non-defective test booklets


1
Expert's answer
2022-02-20T17:24:03-0500

The number of possible outcomes = 23 = 8

Thus the sample space S = { DDD, DDN, DND, DNN, NNN, NND, NDN, NDD}


Define X as the random variable representing the number of non-defective test booklets, then

p(x=0) = 1/8

p(x=1) = 3/8

p(x=2) = 3/8

p(x=3) = 1/8


Hence we may represent the probability distribution as below

x 0 1 2 3

p(x) 1/8 3/8 3/8 1/8